OpenStudy (anonymous):

If a pro basketball player has a vertical leap of about 20 inches, what is his hang time? Use the hang-time function V=48t^2

OpenStudy (dumbcow):

20=48t^2 5/12 = t^2 sqrt(5/12) = t

OpenStudy (anonymous):

So what is the exact hang time of the basketball player in seconds?

OpenStudy (dumbcow):

sqrt(5/12)

OpenStudy (dumbcow):

\[\sqrt{\frac{5}{12}} = \frac{\sqrt{5}}{2\sqrt{3}} = \frac{\sqrt{15}}{6}\]

OpenStudy (anonymous):

Thank you